Flag pole is 24 feet and anchor cable is 10 feet away. How long is anchor cable?
Answer by unlockmath(1688)

Hello,
The shape forms a right triangle. Since we know the two legs are 24 and 10 we can find the hypotenuse (the diagonal line) using the pythagorean equation
a^2 + b^2 = c^2 Plug in the numbers
24^2 + 10^2 = c^2 Gives us:
576 + 100 = c^2
676 = c^2 Square root both sides turns out as:
26 = c There you go. The length of the anchor cable is 26 ft
RJ Toftness
www.math-unlock.com